Assume the nominal rate of return is 5.08% and the inflation rate is 6.10%. Find the real rate of return using the exact formula.

Answer Format: Percentage Round to: 0 decimal places (Example: 9%, % sign required. Will accept decimal format rounded to 2 decimal places (ex: 0.09))
Answer #1

Solution :-

(1 + Real Rate of Return) = (1+Nominal Return)/(1+Inflation)

(1 + Real Rate of Return) = (1+0.0508)/(1+0.061)

(1 + Real Rate of Return ) = 0.99038

Real rate of return = 0.99038 - 1 = -0.0096 = -1%